# Explainable Molecular Property Prediction: Aligning Chemical Concepts with Predictions via Language Models

### v2

@@ Abstract
Providing explainable molecular property predictions is critical for many scientific domains, such as drug discovery and material science. Though transformer-based language models have shown great potential in accurate molecular property prediction, they neither provide chemically meaningful explanations nor faithfully reveal the molecular structure-property relationships. In this work, we develop a new framework for explainable molecular property prediction based on language models, dubbed as Lamole, which can provide chemical concepts-aligned explanations  [^1] Lamole is from the name of a historical winery in Italy called Lamole di Lamole.. We first leverage a designated molecular representation — the Group SELFIES — as it can provide chemically meaningful semantics. Because attention mechanisms in Transformers can inherently capture relationships within the input, we further incorporate the attention weights and gradients together to generate explanations for capturing the functional group interactions. We then carefully craft a marginal loss to explicitly optimize the explanations to be able to align with the chemists’ annotations. We bridge the manifold hypothesis with the elaborated marginal loss to prove that the loss can align the explanations with the tangent space of the data manifold, leading to concept-aligned explanations. Experimental results over six mutagenicity datasets and one hepatotoxicity dataset demonstrate Lamole can achieve comparable classification accuracy and boost the explanation accuracy by up to $14.8\%$, being the state-of-the-art in explainable molecular property prediction.
@@ 1 Introduction
Molecular property prediction aims to reveal the relationships between molecular structures and their properties, assisting scientists in screening molecules for various applications such as drug discovery and material design [1, 2, 3, 4, 5]. Several learning-based models are devised based on the underlying molecular representations, such as graph-based molecular representations and string-based molecular representations. Among them, string-based molecular representations, e.g., simplified molecular input line entry systems (SMILES [6]), stand out for their simplicity and adaptability [2, 7, 8]. By viewing the string-based molecular representation as a form of "chemical" language, the Transformer-based language models (LMs) like Bert [9] offer higher throughput and accuracy for molecular property prediction [2, 10].
Figure 1: (a) The molecule visualization of prediction/explanation. The interaction between the benzene ring and the nitro group (highlighted in red) induces the mutagenic property of the molecule. (b)-(e) are the explanation results obtained with various methods: (b) OrphicX [11]; (c) GNNExplainer [12], (d) GNN with gradient-based explainability technique (GradCAM [13]); (e) Bert with GradCAM (molecular string SMILES as input); (f) Bert with GradCAM (molecular string Group SELFIES [8] as the input representation); (g) Our method Lamole assigns an importance score to each functional group/fragment to indicate their contribution to the property.
Despite the superior performance of learning-based prediction methods, what key factors induce the model’s predictions remain largely unexplored, impeding further advancements in the scientific domains. Typically, it is crucial to obtain explanations of predictions while achieving accurate predictions. These obtained explanations could be used for scientific hypotheses validation or/and providing actionable insights for refining investigations, such as optimization for molecular structural design [14, 15, 16]. With different types of molecular representations, explainability techniques for graph neural networks (GNNs) or LMs might be adopted to alleviate the general lack of explainability in molecular prediction [17, 18, 19, 20].
@@
However, we argue that existing explainability techniques often struggle to generate plausible explanations that can highlight chemically meaningful substructures and faithfully uncover the structure-property relationships simultaneously. Specifically, 1) from the molecular representation perspective, the commonly used representations do not explicitly encode the chemically meaningful substructures; current explainability methods can only highlight individual atoms and bonds as explanations (see Fig. 1 (b)$\sim$(e)). 2) From the perspective of explainability techniques, current methods suffer from two main limitations. First, they cannot effectively capture the interactions between functional groups within the molecular structure. Second, they could not generate explanations that align with the domain knowledge of chemists. As a result, they fail to produce explanations that faithfully reflect the structure-property relationships. (see Fig. 1 (f)). Therefore, a novel framework is imperative to provide plausible explanations for molecule property predictions.
The recently proposed string-based molecular representation — Group SELFIES [8] — encodes molecules at the functional group/fragment level, showcasing the possibility of obtaining chemically meaningful explanations. As shown in Fig. 2, Group SELFIES converts a p-nitrobenzoic acid molecule to a string, which explicitly encodes functional groups as tokens, including a benzene, a nitro group, and a carboxyl group. Therefore, with the rich chemical semantics of Group SELFIES, this work develops a new explainable molecular property prediction framework that can provide chemical concepts-aligned explanations based on language models, called Lamole (see Fig. 1 (g)). In particular, we address the above issues with the following contributions.
@@
    Lamole uses the molecular representation — Group SELFIES — as input tokens for providing the chemically meaningful semantics. By deriving the gradients regarding the input functional group/fragment tokens, gradient-based explanations are generated to indicate the contribution of each functional group/fragment to the molecular properties.
@@
    By disentangling the information flows of Transformer-based LMs, we observe that the gradient-based explanations cannot well capture the functional group interactions. Therefore, we further integrate attention weights into gradients to generate explanations that can perceive the functional group interactions. The intuition is that the attention weights in the information flows can inherently capture the functional group interactions within the input.
@@
Experimental results over six mutagenicity datasets and one hepatotoxicity dataset demonstrate that Lamole can achieve comparable classification accuracy and improve the explanation accuracy by up to $14.8\%$. We also quantitatively evaluate explanations based on the first proposed plausibility metric. Compared to alternative baselines, the explanation plausibility of Lamole is improved by up to $9\%$. Extensive experimental studies demonstrate Lamole achieves state-of-the-art performance in explainable molecular property prediction.
@@ 2 Related Work
Explainable Molecular Property Prediction. To gain insights into the model predictions, several explainable GNNs are proposed to explain the relationship between the input molecular graph and the prediction [11, 12, 21, 20]. However, as shown in Fig. 1 (b)$\sim$(d), they only provide explanations at individual atoms and bond levels instead of chemically meaningful substructures. On the other hand, with string-based molecular representations, LMs show great potential in molecular property prediction [10, 22, 23, 24]. However, the "black-box" characteristics of LMs hamper trust and efficient use of these potent computational tools in scientific domains.
Explainability Techniques in LMs. Transformers, lying the cornerstone of LMs, have advanced the state-of-the-art in various research fields, including molecular property prediction [10, 22, 24]. One way to explain their output is to use the attention weights over the input tokens. However, recent studies suggest that "attention is not explanation" because attention weights are often inconsistent with the feature importance [25, 26, 27]. Perturbation-based methods perturb the inputs and evaluate the output changes to reveal the relationship between the input and output. However, the generated explanations may change drastically with very small perturbations [28]. Gradient-based methods determine the feature importance by the partial derivatives of the output to each feature [13]. However, several works show that the gradient-based methods may not be reliable, as they disregard the influence of model architectures on the output and fail to incorporate the information of the model architectures into the explanations [29, 28, 30]. To alleviate this issue, by disentangling the model architectures’ information flows, we combine the attention weights and gradients to generate explanations that faithfully reveal the structure-property relationship.
@@ 3 Methodology
Problem Setup. Given a molecular graph $g^{(i)}\in\mathcal{G}=\{(g^{(i)},y^{(i)})\}$, explainable molecular property prediction aims to train a model $f$ to map a molecule $g$ to its property $y$, denoted as $f:g\rightarrow y$, while providing an importance score vector $\mathbf{v}^{(i)}=\{\mathbf{v}_{1}^{(i)},\ldots,\mathbf{v}_{j}^{(i)}\}$ to indicate the contribution of $j$-th functional group/fragment to its property $y$. Particularly, this work proposes to convert the molecular graph $g^{(i)}$ into the Group SELFIES string represented as $s^{(i)}=\{t_{1}^{(i)},\ldots,t_{j}^{(i)}\}$, where $t_{j}^{(i)}$ is the $j$-th functional group/fragment’s token, i.e., [$\cdot$] in Fig. 2. In addition, this work calibrates the explanations in a supervised manner. For this purpose, a few molecules with annotation masks are provided. The annotation masks $\mathbf{m}(g^{(i)})\in\{0,1\}^{j}$ indicate whether $t_{j}^{(i)}$ is the token of a ground truth substructure, where $\mathbf{m}_{j}(g^{(i)})=1$ denotes the substructure corresponding to the token $t_{j}^{(i)}$ inducing the molecular property of $g^{(i)}$. This work uses $\mathcal{D}=\{(s^{(i)},y^{(i)})\}$ with a few annotation masks to learn a model $f$ for explainable molecular property prediction. We omit the superscript (i) for simplicity in the following parts.
@@ 3.1 Our Design: Lamole
In this work, we use pre-trained Transformer-based LMs, e.g., Bert family models, as our backbone. Then, we fine-tune the LMs on the Group SELFIES corpus for explainable molecular property prediction. An illustration of proposed Lamole is in Fig. 3. In what follows, we will introduce the detailed design of Lamole.
@@
Information Flow-Based Explanations. To derive the explanations from Transformer-based LMs, a common practice is to use gradient-based methods to determine each feature’s importance by analyzing the output’s partial derivatives to each input feature [13]. However, we argue that using gradients alone cannot effectively capture functional group interactions. As depicted in Fig. 1 (f), the gradient-based method CradCAM incorrectly attributes the property to the nitro group and a benzene ring that is not connected to the nitro group.
@@
Firstly, we show the process of deriving the gradient-based explanations. Similar to GradCAM [13], the gradient with respect to the $j$-th token’s embedding $\mathbf{h}_{j}^{(l)}$ at the $l$-th layer is derived by
$\nabla\mathbf{h}_{j}^{(l)}=\frac{\partial\hat{y}}{\partial\mathbf{h}_{j}^{(l)}},$ (1)
where $\nabla\mathbf{h}_{j}$ signifies the importance of the $j$-th token in relation to the predicted property $\hat{y}$. Due to the skip connection in Fig. 3 right panel, the input and its corresponding gradient should be leveraged together, and the weighted importance $\mathbf{w}$ of the $j$-th token at the $l$-th layer can be determined by
@@
First, we formally define the plausibility of explanations. "Plausibility" refers to how the interpretation convinces humans [31, 32, 33]. Similarly, in our context, "plausibility" refers to the degree of confidence in the explanations that would convince the chemists.
@@
In our work, the ground truth substructures are annotated by a binary mask vector $\mathbf{m}(g)\in\{0,1\}^{j}$. It is worth noting that using only a few annotations can significantly improve the explanation accuracy. Specifically, the mask vector $\mathbf{m}$ enforces the explanations to align with the ground truth substructures. To achieve the goal, a max-margin loss is designed by maximizing the mean value of the importance scores of tokens that have mask values of 1 while minimizing the mean value of importance scores for tokens with mask values of 0.
@@
where $\triangle_{1}$ is a margin term, $N_{s}$ is the number of tokens with mask values $\mathbf{m}(g)$ of 0, and $N_{c}$ is the number of tokens with mask values $\mathbf{m}(g)$ of 1. The overall optimization objective of the fine-tuning stage is to minimize $\mathcal{L}_{CE}+\mathcal{L}_{M}$. By minimizing $\mathcal{L}_{M}$, Lamole aligns the explanations with the domain knowledge, enhancing the plausibility of explanations.
The next section will theoretically show that by using the designed marginal loss, the explanations can faithfully reflect the structure-property relationships.
@@ 3.2 Theoretical Analysis
Manifold Hypothesis. It is widely believed that natural data, including molecules, distribute around a manifold [34, 11, 35, 36]. According to the manifold hypothesis for gradient-based explanations [34], if a feature lies in the tangent space of a manifold, then the feature respects the manifold and contributes to the class, and such a feature is desirable to be explained. We call these features "causal features" in our work. Conversely, if a feature is orthogonal to the manifold, then the feature does not contribute to the class. We call these features "spurious features".
@@
With this decomposition, we demonstrate how the gradient-based explanations $\nabla_{g}\log p(y|g)$ can uncover the structure-property relationships. Due to the page limitation, we provide the proof in Appendix 1.

### v3

@@ Abstract
Providing explainable molecular property predictions is critical for many scientific domains, such as drug discovery and material science. Though transformer-based language models have shown great potential in accurate molecular property prediction, they neither provide chemically meaningful explanations nor faithfully reveal the molecular structure-property relationships. In this work, we develop a framework for explainable molecular property prediction based on language models, dubbed as Lamole, which can provide chemical concepts-aligned explanations  [^1] Lamole is from the name of a historical winery in Italy called Lamole di Lamole.. We take a string-based molecular representation — Group SELFIES — as input tokens to pre-train and fine-tune our Lamole, as it provides chemically meaningful semantics. By disentangling the information flows of Lamole, we propose considering both self-attention weights and gradients for better quantification of each chemically meaningful substructure’s impact on the model’s output. To make the explanations more faithfully respect the structure-property relationship, we then carefully craft a marginal loss to explicitly optimize the explanations to be able to align with the chemists’ annotations. We bridge the manifold hypothesis with the elaborated marginal loss to prove that the loss can align the explanations with the tangent space of the data manifold, leading to concept-aligned explanations. Experimental results over six mutagenicity datasets and one hepatotoxicity dataset demonstrate Lamole can achieve comparable classification accuracy and boost the explanation accuracy by up to $14.3\%$, being the state-of-the-art in explainable molecular property prediction.
@@ 1 Introduction
Molecular property prediction aims to reveal the molecular structures-property relationships, assisting scientists in screening molecules for various applications such as drug discovery and material design (Fang et al., 2022; Deng et al., 2023; Tripp et al., 2023; Wang et al., 2024; Ekström Kelvinius et al., 2023; Hong et al., 2024). Several learning-based models are devised based on the underlying molecular representations, such as graph-based and string-based molecular representations. Among them, string-based molecular representations, e.g., simplified molecular input line entry systems (SMILES (Weininger, 1988)), stand out for their simplicity and adaptability (Deng et al., 2023; Wigh et al., 2022; Cheng et al., 2023). By viewing the string-based molecular representation as a form of "chemical" language, the Transformer-based language models (LMs) like Bert (Kenton & Toutanova, 2019) offer higher throughput and accuracy for molecular property prediction (Deng et al., 2023; Chithrananda et al., 2020).
Figure 1: (a) The molecule visualization of prediction/explanation. The interaction between the benzene ring and the nitro group (highlighted in red) induces the mutagenic property of the molecule. (b)-(e) are the explanation results obtained with various methods: (b) OrphicX (Lin et al., 2022); (c) GNNExplainer (Ying et al., 2019), (d) GNN with gradient-based explainability technique (GradCAM (Selvaraju et al., 2017)); (e) Bert with GradCAM (molecular string SMILES as input); (f) Bert with GradCAM (molecular string Group SELFIES (Cheng et al., 2023) as the input representation); (g) Our method Lamole assigns an importance score to each functional group/fragment to indicate their contribution to the property.
@@
Despite the superior performance of learning-based prediction methods, what key factors induce the model’s predictions remain largely unexplored, impeding further advancements in the scientific domains. Typically, it is crucial to obtain explanations of predictions while achieving accurate predictions. These obtained explanations could be used for scientific hypotheses validation or/and providing actionable insights for refining investigations, such as optimization for molecular structural design (Wu et al., 2023; Wellawatte et al., 2023; Das et al., 2022). With different types of molecular representations, explainability techniques for graph neural networks (GNNs) or LMs might be adopted to alleviate the general lack of explainability in molecular prediction (Proietti et al., 2024; Ying et al., 2019; Ye et al., 2023; Lin et al., 2021).
However, we argue that existing explainability techniques often struggle to generate plausible explanations that can highlight chemically meaningful substructures and faithfully uncover the structure-property relationships simultaneously. Specifically, 1) from the molecular representation perspective, the commonly used representations do not explicitly encode the chemically meaningful substructures; current explainability methods can only highlight individual atoms and bonds as explanations (see Fig. 1 (b)$\sim$(e)). 2) From the perspective of explainability techniques, current methods suffer from two main limitations. First, they cannot effectively capture the interactions between functional groups within the molecular structure. Second, they could not generate explanations that align with chemists’ intuition. As a result, they fail to produce explanations that faithfully reflect the structure-property relationships. (see Fig. 1 (f)). Therefore, an effective framework is imperative for explainable molecule property predictions.
The recently proposed string-based molecular representation — Group SELFIES (Cheng et al., 2023) — encodes molecules at the functional group/fragment level, showcasing the possibility of obtaining chemically meaningful explanations. As shown in Fig. 2, Group SELFIES converts a p-nitrobenzoic acid molecule to a string, which explicitly encodes chemically meaningful substructures as tokens, including a benzene, a nitro group, and a carboxyl group. Compared with 2D molecular graphs, Group SELFIES provides inherent semantic information, making it easier for the model to capture and understand chemically meaningful semantics. Moreover, using Group SELFIES eliminates the need to identify or segment chemically meaningful substructures in 2D molecular graphs. With Group SELFIES’s simplicity and adaptability, this work develops an explainable molecular property prediction framework based on language models to provide chemical concepts-aligned explanations, called Lamole (see Fig. 1 (g)). The contributions can be summarized as follows.
@@
    We found that existing explainability techniques fail to provide chemically meaningful explanations, perceive functional group interactions, and reveal molecular structure-property relationships faithfully. To address the issues, we use Group SELFIES to pre-train and fine-tune LMs to make LMs easily understand the chemically meaningful semantics. Moreover, the process of generating explanations should reflect the reasoning process behind the model architecture. Therefore, by disentangling the information flows of Transformer-based LMs, we integrate the self-attention weights and gradients to capture the substructure interactions to better quantify each chemically meaningful substructure’s contribution to the predicted molecular properties.
@@
Experimental results over six mutagenicity datasets and one hepatotoxicity dataset demonstrate that Lamole can achieve comparable classification accuracy and improve the explanation accuracy by up to $14.3\%$. We also quantitatively evaluate explanations based on the first proposed plausibility metric. Compared to alternative baselines, the explanation plausibility of Lamole is improved by up to $9\%$. Extensive experimental studies demonstrate Lamole achieves state-of-the-art performance in explainable molecular property prediction.
@@ 2 Related Work
Several explainable GNNs are proposed to explain the relationship between the input graph and the prediction (Sun et al., 2023; Xiong et al., 2019; Lin et al., 2022; Ying et al., 2019; Luo et al., 2020; Lin et al., 2021). Among these works, structure similarity or attention weights are proposed to capture structural interaction (Sun et al., 2023; Xiong et al., 2019). However, two similar substructures do not necessarily lead to interaction between them, and attention weights are often inconsistent with the feature importance (Jain & Wallace, 2019; Serrano & Smith, 2019; Abnar & Zuidema, 2020). In addition, as shown in Fig. 1 (b)$\sim$(d), some trivial structures received relatively high importance scores, indicating the explanations might not align well with the chemical concepts.
On the other hand, with string-based molecular representations, LMs show great potential in molecular property prediction (Chithrananda et al., 2020; Wang et al., 2019; Ahmad et al., 2022; Ross et al., 2022). However, the "black-box" characteristics of LMs hamper trust use of these potent computational tools in scientific domains. Some explainability techniques could be applied to LMs. One way is to use the attention weights over the input tokens. However, recent studies suggest that "attention is not explanation" because attention weights could not reflect the true feature importance (Jain & Wallace, 2019; Serrano & Smith, 2019; Abnar & Zuidema, 2020). Perturbation-based methods perturb the inputs and evaluate the output changes to reveal the input importance. However, the generated explanations may change drastically with very small perturbations (Agarwal et al., 2021). Gradient-based methods determine the feature importance by the partial derivatives of the output to each feature (Selvaraju et al., 2017). However, several works show that the gradient-based methods may not be reliable, as they disregard the influence of model architectures on the output and fail to incorporate the information of the model architectures into the explanations (Adebayo et al., 2018; Agarwal et al., 2021; Rudin, 2019). Therefore, the explanation generation process should reflect the model reasoning process behind the model architectures. To this end, we disentangle the model architectures’ information flows to generate explanations that faithfully reveal the structure-property relationship.
@@ 3 Methodology
Problem Setup. Given a dataset $\mathcal{G}=\{(g^{(i)},y^{(i)})\}$ consisting of molecular graphs $\{g^{(i)}\}$ with their property labels $\{y^{(i)}\}$, explainable molecular property prediction aims to train a model $f$ to map a molecule $g$ to its property $y$, denoted as denoted as $f:g\mapsto y$, while providing an importance score vector $\mathbf{v}^{(i)}=\{\mathbf{v}_{1}^{(i)},\ldots,\mathbf{v}_{j}^{(i)}\}$ to indicate the contribution of $j$-th functional group/fragment to its property $y$. Particularly, this work proposes to convert the molecular graph $g^{(i)}$ into the Group SELFIES string represented as $s^{(i)}=\{t_{1}^{(i)},\ldots,t_{j}^{(i)}\}$, where $t_{j}^{(i)}$ is the $j$-th functional group/fragment’s token, i.e., [$\cdot$] in Fig. 2. In addition, this work calibrates the explanations in a supervised manner. For this purpose, a few molecules with annotation masks are provided. The annotation masks $\mathbf{m}(g^{(i)})\in\{0,1\}^{j}$ indicate whether $t_{j}^{(i)}$ is the token of a ground truth substructure, where $\mathbf{m}_{j}(g^{(i)})=1$ denotes the substructure corresponding to the token $t_{j}^{(i)}$ inducing the molecular property of $g^{(i)}$. This work uses $\mathcal{D}=\{(s^{(i)},y^{(i)})\}$ with a few annotation masks to learn a model $f$ for explainable molecular property prediction. We omit the superscript (i) for simplicity in the following parts.
@@ 3.1 Our Design: Lamole
In this work, we pre-train Transformer-based LMs, e.g., Bert family models, using the Group SELFIES corpus to make the models understand the chemical semantics behind Group SELFIES strings. Then, we fine-tune the LMs with Group SELFIES strings and the molecular property labels for explainable molecular property prediction. An illustration of proposed Lamole is in Fig. 3. In what follows, we will introduce the detailed design of Lamole.
@@
Information Flow-Based Explanations. To derive the explanations from Transformer-based LMs, a common practice is to use gradient-based methods to determine each feature’s importance by analyzing the output’s partial derivatives to each input feature (Selvaraju et al., 2017). However, we argue that using gradients alone cannot effectively capture substructure interactions. As depicted in Fig. 1 (f), the gradient-based method CradCAM incorrectly attributes the property to the nitro group and a benzene ring that is not connected to the nitro group.
@@
Firstly, we show the process of deriving the gradient-based explanations. Similar to GradCAM (Selvaraju et al., 2017), the gradient with respect to the $j$-th token’s embedding $\mathbf{h}_{j}^{(l)}$ at the $l$-th layer is derived by $\nabla\mathbf{h}_{j}^{(l)}=\partial\hat{y}/{\partial\mathbf{h}_{j}^{(l)}}$, where $\nabla\mathbf{h}_{j}$ signifies the importance of the $j$-th token in relation to the predicted property $\hat{y}$. Due to the skip connection in Fig. 3 right panel, the input, and its corresponding gradient should be leveraged together, and the weighted importance $\mathbf{w}$ of the $j$-th token at the $l$-th layer can be determined by
@@
First, we formally define the plausibility of explanations. "Plausibility" refers to how the interpretation convinces humans (Wiegreffe & Pinter, 2019; Herman, 2017; Jacovi & Goldberg, 2020). Similarly, in our context, "plausibility" refers to the degree of confidence in the explanations that would convince the chemists.
@@
Eq. (4) defines the explanation plausibility based on the scores of two parts, i.e., the scores on ground truth and the scores on non-ground truth. The lower the scores on non-ground truth and the greater the scores on ground truth, the better explanation plausibility. Therefore, to maximize the plausibility, our objective can be transformed to minimize the importance score of non-ground truth and maximize the importance score of ground truth.
To this end, we design a max-margin loss to optimize the importance score. In our work, the ground truth substructures are annotated by a binary mask vector $\mathbf{m}(g)\in\{0,1\}^{j}$. It is worth noting that using only a few annotations can significantly improve the explanation accuracy. Specifically, the mask vector $\mathbf{m}$ enforces the explanations to align with the ground truth substructures. To achieve the goal, a max-margin loss is designed by maximizing the mean value of the importance scores of tokens that have mask values of 1 while minimizing the mean value of importance scores for tokens with mask values of 0.
@@
where $\triangle_{1}$ is a margin term, $N_{s}$ is the number of tokens with mask values $\mathbf{m}(g)$ of 0, and $N_{c}$ is the number of tokens with mask values $\mathbf{m}(g)$ of 1. The overall optimization objective of the fine-tuning stage is to minimize $\mathcal{L}_{CE}+\mathcal{L}_{M}$. The core of Eq. (5) is the discrepancy between the average importance score of ground truth and the average importance score of non-ground truth. By minimizing $\mathcal{L}_{M}$, the discrepancy between the two importance scores is maximized. In other words, the average importance score of non-ground truth is suppressed, and the average importance score of ground truth is increased. Finally, the explanation plausibility defined in Eq. (4) is improved. The next section will theoretically show that by using the designed marginal loss, the explanations can faithfully reflect the structure-property relationships.
@@ 3.2 Theoretical Analysis
Manifold Hypothesis. It is widely believed that natural data, including molecules, distribute around a manifold (Bordt et al., 2023; Lin et al., 2022; Godwin et al., 2022; Singh et al., 2020). According to the manifold hypothesis for gradient-based explanations (Bordt et al., 2023), if a feature lies in the tangent space of a manifold, then the feature respects the manifold and contributes to the class, and such a feature is desirable to be explained. We call these features "causal features" in our work. Conversely, if a feature is orthogonal to the manifold, then the feature does not contribute to the class. We call these features "spurious features".
@@
With this decomposition, we demonstrate how the gradient-based explanations $\nabla_{g}\log p(y|g)$ can uncover the structure-property relationships. Due to the page limitation, we provide the proof in Appendix.